OpenStudy (anonymous):

Use the function defined below to find an equation of the tangent line at the point indicated. y = cos(x), x = pi/3

OpenStudy (anonymous):

I got the slope of \[-\sqrt{3}/2\] I am having trouble making the equation

OpenStudy (anonymous):

do you have the point?

OpenStudy (anonymous):

x=pi/3 y=cos(pi/3)=1/2

OpenStudy (anonymous):

it is \[(\frac{\pi}{3},\frac{1}{2})\]so point - slope it

OpenStudy (anonymous):

m=-(sqrt3)/2

myininaya (myininaya):

\[(x=\frac{\pi}{3},y=?)\]

OpenStudy (anonymous):

plug the values in the formula

OpenStudy (anonymous):

I need step by step AFTER you get the slope.

OpenStudy (anonymous):

y-(1/2)=(-sqrt3)/2(x-1/2)

myininaya (myininaya):

uzma is that right?

myininaya (myininaya):

i think it should be y-y_1=m(x-x_1)

OpenStudy (anonymous):

isnt it?

OpenStudy (anonymous):

yes right y1=1/2 x1=pi/3

OpenStudy (anonymous):

\[m=-\sqrt{3}/2\]

myininaya (myininaya):

\[y-\frac{1}{2}=- \frac{\sqrt{3}}{2}(x-?)\]

OpenStudy (anonymous):

I'm so lost.

OpenStudy (anonymous):

yes it should be x-pi/3

myininaya (myininaya):

right he just used point slope formula

OpenStudy (anonymous):

\[y-\frac{1}{2}=-\frac{\sqrt{3}}{2}(x-\frac{\pi}{3})\] is my answer

OpenStudy (anonymous):

Look, I have the slope. Can somebody PLEASE give me STEP BY STEP explanation of what the heck I need to do.

myininaya (myininaya):

we agree with you satellite

OpenStudy (anonymous):

then with slope either you should have a y intercept or a point on the line

myininaya (myininaya):

ham if you plug x into get y there we know a point on the line that is the point we plugged in into the point slope formula

OpenStudy (anonymous):

yes you use \[y-y_1=m(x-x_1)\] with \[m=-\frac{\sqrt{3}}{2}, x_1=\frac{\pi}{3}, y_1=\frac{1}{2}\]

myininaya (myininaya):

my sentences suck

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you r given x=pi/2, using this in the curve u can find y which is -(sqrt3)/2

OpenStudy (anonymous):

Why didn't you put it into y=mx+b form?

myininaya (myininaya):

you can

OpenStudy (anonymous):

what is b over here?

myininaya (myininaya):

\[y=\cos(x)=\frac{1}{2} when x=\frac{\pi}{3}\] \[y=mx+b\] \[\frac{1}{2}=\frac{-\sqrt{3}}{2}\cdot \frac{\pi}{3}+b\] solve for b the y-intercept \[\frac{1}{2}+\frac{\sqrt{3}\pi}{6}=b\] so we can write: \[y=\frac{-\sqrt{3}}{2}x+\frac{1}{2}+\frac{\sqrt{3} \pi}{6}\]